Please show work as to how you got your answers. For calculator solutions, please give calculator inputs.

You have the following details about the First Banc Corp. bond: Maturity = 9 years, Coupon rate (annual) = 5.25%, Yield to maturity = 4.2%, Par value = $1,000. Calculate the value of the bond, its duration and modified duration. What will be the change in value if interest rates go up by 1.0%?
Answer #1

All the calculations and formulas used are explained in the below images in detail.

1) Value of the bond = $1000.077

Sol & Giren, Maturity = 9years. Coupon rate & 5.250%., t. YTM= 4.2% 2 Parvalue = $1000. I . a) Value at the bondir Vo= l + TP & value of * 7) The bond (14 )9 - 1000.017$.2) Duration of the bond = 7.4 years.

(b) Duration of the bonds Macauley at - exel (4 *) nx (1 M ) Duration I*(-15) + Maison Ix/1- +3) When This is I = 52.5 ; n =.. Duration = 1,719.901 + 9x 1000 . (1+0.0525) 52-5*(-CHO0259 ) + 1000 0.0525 / (1+0.0525) 9 21,919.901 € 5,678.63 + 630.96 73) Modified duration of the bond = 7.1 years

19 on lali Tion Modified ducation of the bond:- Macauley Modified duration - Duration (HKTM) When, n= No.of Coupons periods p4) Value of the bond when interest rates are increased by 1%

(d) value of the bands I = $10000 6.25% = 62.56 son klhen, it increases by 1% 9. e., 6.25%. 1 year I ((1+raja) = (1 +0.0625)